# Resonant absorption and linear photovoltaic effect in ferroelectric moiré heterostructures

### v1

@@ Abstract
Since an account of finite out-of-plane electric field does not change results obtained for zero electric field, in the main text we focus on the latter case. The corresponding results for a finite field are gathered in sections S4 and S5 of SM. In the absence of the out-of-plane electric field the moiré potential (1) possesses odd parity, $V(-\bm{r})=-V(\bm{r})$, illustrated on the bottom insets in Fig. S5 ($V^{(s)}_{l}\equiv 0$, $l=0,1,2,\dots$ ). The operator $\hat{I}$, transforming $\bm{r}\to-\bm{r}$, then, corresponds to the chiral symmetry of Hamiltonian (3), satisfying $\hat{H}\hat{I}=-\hat{I}\hat{H}$. Combined with an intravalley hidden symmetry $\hat{T}=i\sigma_{y}\hat{K}$ (where $\hat{K}$ is the complex conjugation operator and $\hat{H}\hat{T}=\hat{T}\hat{H}$), this leads to electron-hole symmetry of states with the same quasi-momentum: $\varepsilon_{n}(\bm{k})=-\varepsilon_{-n}(\bm{k})$. We also note that the Hamiltonian (3) is invariant under non-centrosymmetric $C_{3v}$ point group, which includes, in particular, mirror reflection with respect to the $y$-axis, represented by an operator, $\hat{M}_{y}=\sigma_{x}\{y\to-y\}$ used below.
In Fig. S5 we demonstrate the dispersions obtained for several values of graphene doping, calculated with the help of expansion of $u_{n,\bm{k}}(\bm{r})$ in Fourier series and diagonalization of the resulting matrix for all quasimomenta $\bm{k}$ in mBZ, with the number of harmonics large enough to satisfy convergence in the low-energy spectral interval. The lowest mini-bands ($n=\pm 1$) are characterised by gapless spectrum at the neutrality point ($\varepsilon=0$) as the electrostatic moiré potential does not break the sublattice symmetry required to produce a gap for mDF. However, the group velocity of $n=\pm 1$ minibands can be substantially suppressed by the moiré potential. To demonstrate this we project out Eq. (1) to a subspace of the two degenerate states $\psi_{\pm 1,\bm{k}=0}$ and perturbatively account for $v\,\bm{\sigma}\cdot\bm{k}$. Then, leaving in Fourier series for $u_{\pm 1,\bm{k}=0}$ only the zeroth and $\pm\bm{g}^{(1)}_{1,2,3}$ harmonics, we obtain the following effective Hamiltonian (see section S2 in SM):
@@
From Eq. (5) it follows that $v_{*}$ is a funciton of both moiré superlattice period ($\approx a/\theta$) and, implicitly, free carrier density via $V_{1}^{(a)}$. Unlike graphene/hBN moiré superlattices [37, 38] or periodically patterned graphene [39, 40] where renormalization of the primary mDF velocity is small, here, decrease of $v_{*}$ can be as large as $\approx 3$ times for realistic parameters (see bottom inset in Fig. S5). Opportunity to tune the mDF velocity can have practical significance for adjusting transmission probability of p-n junctions [41] in the heterostructures, whereas its reduction may help to observe Landau level breakdown even at moderate homogeneous electric fields [42, 43].
@@
where +/+/- sign is for electrons/holes, $\pm\varepsilon_{\bm{\mu}}=\pm\sqrt{(\hbar v_{*}g)^{2}+\left(V_{1}^{(a)}\right)^{2}}/2$ are energies of the secondary Dirac points ($g=|\bm{g}_{1}^{1}|$), and $\sigma_{0}$ is an identity matrix. Hamiltonian (6) gives rise to strongly anisotropic mDF (see Fig. S5), where in contrast to graphene/hBN [37, 38, 45] or patterned graphene [39] velocities in both principal axes can be controlled by carrier concentration and twist angle.
@@
Figure 2: Frequency dependences of the absorption coefficient (7) for doping levels corresponding to Fig. S5 at $T\ll\hbar\omega_{3}$. The resonances, $\omega_{1,2,3}$, corresponds to energy differences between states at peaks of DOS, shown in Fig. S5, where the interband velocity matrix elements are non-zero. Dashed line indicates absorption ($2.3\%$) of an isolated graphene layer [47].
The $\alpha(\omega)$ dependences, displayed in Fig. 2 for $\theta=0.18^{\circ}$, exhibit resonances whose frequencies depend on the filling of mini-band structure. At an electron concentration of $n=10^{10}$ cm-2 (corresponding to filling of $n=1$ mini-band before the Lifshitz transition, left panel in Fig. S5), the absorption coefficient is dominated by three terms in Eq. (7): $(n^{\prime},n)=(-1,1)$, (1,2)(-1,2) and (2,1)(-2,1). The first term accounts for the peak at $\omega_{1}$, while the latter two contribute equally to the peak at $\omega_{2}$.
At a higher concentration, $n=10^{11}$ cm-2 (middle panel in Fig. S5), transitions between electron and hole mini-bands are suppressed by Pauli blocking, whereas transition between electron $n^{\prime}=1$ and $n=2$ mini-bands become activated in Eq. (7), resulting in a peak at frequency $\omega_{3}$. In both cases, the resonances arise from allowed optical transitions from filled states at one vHS to empty states of another vHS forming close to the boundary of mid- and far-infrared range.

### v2

@@ Abstract
Since an account of finite out-of-plane electric field does not change results obtained for zero electric field, in the main text we focus on the latter case. The corresponding results for a finite field are gathered in sections S4 and S5 of SM. In the absence of the out-of-plane electric field the moiré potential (1) possesses odd parity, $V(-\bm{r})=-V(\bm{r})$, illustrated on the bottom insets in Fig. 1 ($V^{(s)}_{l}\equiv 0$, $l=0,1,2,\dots$ ). The operator $\hat{I}$, transforming $\bm{r}\to-\bm{r}$, then, corresponds to the chiral symmetry of Hamiltonian (3), satisfying $\hat{H}\hat{I}=-\hat{I}\hat{H}$. Combined with an intravalley hidden symmetry $\hat{T}=i\sigma_{y}\hat{K}$ (where $\hat{K}$ is the complex conjugation operator and $\hat{H}\hat{T}=\hat{T}\hat{H}$), this leads to electron-hole symmetry of states with the same quasi-momentum: $\varepsilon_{n}(\bm{k})=-\varepsilon_{-n}(\bm{k})$. We also note that the Hamiltonian (3) is invariant under non-centrosymmetric $C_{3v}$ point group, which includes, in particular, mirror reflection with respect to the $y$-axis, represented by an operator, $\hat{M}_{y}=\sigma_{x}\{y\to-y\}$ used below.
In Fig. 1 we demonstrate the dispersions obtained for several values of graphene doping, calculated with the help of expansion of $u_{n,\bm{k}}(\bm{r})$ in Fourier series and diagonalization of the resulting matrix for all quasimomenta $\bm{k}$ in mBZ, with the number of harmonics large enough to satisfy convergence in the low-energy spectral interval. The lowest mini-bands ($n=\pm 1$) are characterised by gapless spectrum at the neutrality point ($\varepsilon=0$) as the electrostatic moiré potential does not break the sublattice symmetry required to produce a gap for mDF. However, the group velocity of $n=\pm 1$ minibands can be substantially suppressed by the moiré potential. To demonstrate this we project out Eq. (1) to a subspace of the two degenerate states $\psi_{\pm 1,\bm{k}=0}$ and perturbatively account for $v\,\bm{\sigma}\cdot\bm{k}$. Then, leaving in Fourier series for $u_{\pm 1,\bm{k}=0}$ only the zeroth and $\pm\bm{g}^{(1)}_{1,2,3}$ harmonics, we obtain the following effective Hamiltonian (see section S2 in SM):
@@
From Eq. (5) it follows that $v_{*}$ is a funciton of both moiré superlattice period ($\approx a/\theta$) and, implicitly, free carrier density via $V_{1}^{(a)}$. Unlike graphene/hBN moiré superlattices [37, 38] or periodically patterned graphene [39, 40] where renormalization of the primary mDF velocity is small, here, decrease of $v_{*}$ can be as large as $\approx 3$ times for realistic parameters (see bottom inset in Fig. 1). Opportunity to tune the mDF velocity can have practical significance for adjusting transmission probability of p-n junctions [41] in the heterostructures, whereas its reduction may help to observe Landau level breakdown even at moderate homogeneous electric fields [42, 43].
@@
where +/+/- sign is for electrons/holes, $\pm\varepsilon_{\bm{\mu}}=\pm\sqrt{(\hbar v_{*}g)^{2}+\left(V_{1}^{(a)}\right)^{2}}/2$ are energies of the secondary Dirac points ($g=|\bm{g}_{1}^{1}|$), and $\sigma_{0}$ is an identity matrix. Hamiltonian (6) gives rise to strongly anisotropic mDF (see Fig. 1), where in contrast to graphene/hBN [37, 38, 45] or patterned graphene [39] velocities in both principal axes can be controlled by carrier concentration and twist angle.
@@
Figure 2: Frequency dependences of the absorption coefficient (7) for doping levels corresponding to Fig. 1 at $T\ll\hbar\omega_{3}$. The resonances, $\omega_{1,2,3}$, corresponds to energy differences between states at peaks of DOS, shown in Fig. 1, where the interband velocity matrix elements are non-zero. Dashed line indicates absorption ($2.3\%$) of an isolated graphene layer [47].
The $\alpha(\omega)$ dependences, displayed in Fig. 2 for $\theta=0.18^{\circ}$, exhibit resonances whose frequencies depend on the filling of mini-band structure. At an electron concentration of $n=10^{10}$ cm-2 (corresponding to filling of $n=1$ mini-band before the Lifshitz transition, left panel in Fig. 1), the absorption coefficient is dominated by three terms in Eq. (7): $(n^{\prime},n)=(-1,1)$, (1,2)(-1,2) and (2,1)(-2,1). The first term accounts for the peak at $\omega_{1}$, while the latter two contribute equally to the peak at $\omega_{2}$.
At a higher concentration, $n=10^{11}$ cm-2 (middle panel in Fig. 1), transitions between electron and hole mini-bands are suppressed by Pauli blocking, whereas transition between electron $n^{\prime}=1$ and $n=2$ mini-bands become activated in Eq. (7), resulting in a peak at frequency $\omega_{3}$. In both cases, the resonances arise from allowed optical transitions from filled states at one vHS to empty states of another vHS forming close to the boundary of mid- and far-infrared range.